NEED ASAP IN 10 MIN PLEASE!!!! THANK YOU!!5 Consider this simplified balance sheet for Geomorph Trading: 170 Current assets $ Long-term assets 570 Current liabilities
Answer #1

(a ): Debt- Equity Ratio= Total liabilities/Equity

= (95+245+140)/260 = 480/260 = 1.85

(b ): Ratio of long term debt to long term equity = 245/260 = 0.94

(c ): Net working capital= Current assets- Current liabilities =170-95 = $75

(d ): Current Ratio= Current assets/Current liabilities = 170/75 = 1.79